OpenStudy (anonymous):

Find the polynomial function with the degree of 3 and the zeros 4i and 5.

OpenStudy (anonymous):

Degree = 3, x= 5, x=4i (x+1)(x-5)(x-4i)(x+4i) Not sure if thats correct...

OpenStudy (anonymous):

x^2-4x-5

OpenStudy (anonymous):

and thats where I get confused

OpenStudy (anonymous):

The zeroes are 5 and 4i... if you have 4i you have -4i y = (x - 5)(x - 4i)(x + 4i) Don't know where you go the (x + 1) from

OpenStudy (anonymous):

y = (x - 5)(x^2 + 16) y = x^3 - 5x^2 + 16x - 80

OpenStudy (anonymous):

ok , so why is it a +16 and not -16 I also get confused on the middle part of the equation is it: (x-5)(x^2+4xi-4xi-16), and since the 4xi cancel each other other thats why we dont use them?

OpenStudy (anonymous):

-4i * 4i = -16i^2 by definition i^2 = -1 so -16 * -1 = +16

OpenStudy (anonymous):

ok, I'll to remember that for the test. Kinda of confusing but I'll get it.